OpenStudy (anonymous):

Use the “difference of squares” rule to factor the following expression:25-9z^2 A.(5-3z)(5-3z) B.-(3z-5) c.25(1-9y^2) D.(5-3z)(5+3z)

OpenStudy (anonymous):

D IS THE COORECT ANSWER

jimthompson5910 (jim_thompson5910):

25-9z^2 = (5-3z)(5+3z) by the difference of squares rule

OpenStudy (anonymous):

CHECK (5-3z)(3+3z) =25+3z-3z-9z^2 =25-9z^2

OpenStudy (anonymous):

so its c

jimthompson5910 (jim_thompson5910):

D

OpenStudy (anonymous):

k